DIFFERENTIAL GEOMETRY/PDE SEMINAR

Thursday, February 5, 2004 Padelford C-36 1:30 pm Unique continuation and inverse problems in anisotropic elasticity

Jenn-Nan Wang (National Taiwan University)

In this talk I would like to discuss the unique continuation property (UCP) for a large class of second order systems, including the anisotropic elasticity system. I will describe applications of the UCP to inverse problems for anisotropic elastic materials. The results presented in this lecture are joint work with G. Nakamura and G. Uhlmann.

For more information about this seminar, visit the DG/PDE Seminar Web page (from the Math Department home page, www.math.washington.edu, follow the link Seminars, Colloquia, and Conferences).

The University of Washington is committed to providing access, equal opportunity and reasonable accommodation in its services, programs, activities, education and employment for individuals with disabilities. To request disability accommodation contact the Disability Services Office at least ten days in advance at: 206-543-6450/V, 206-543-6452/TTY, 206-685-7264 (FAX), or dso@u.washington.edu.